### v2

# A Note on a Recent Attempt to Prove the Irrationality of $\zeta(5)$

Abstract.
Recently, Shekhar Suman [5] made an attempt to prove the irrationality of $\zeta(5)$. Unfortunately, the proof is not correct. In this note, we discuss the fallacy in the proof.
2020 Mathematics Subject Classification:
Primary 11J72; Secondary 11M06
Keywords: Irrationality, Odd zeta values
1. Introduction
The study of irrationality and transcendence of numbers is a classical topic in transcendental number theory. Recall the Riemann zeta values
$\zeta(n)=\sum_{k\geq 1}\frac{1}{k^{n}}\quad\text{for }n\geq 2.$
Euler proved that
$\zeta(2)=\frac{\pi^{2}}{6}$
and more generally
$\zeta(2n)=-\frac{B_{2n}(2\pi i)^{2n}}{2(2n)!}\quad\text{for }n\geq 1,$
where $B_{m}$ is the $m$-th Bernoulli number. It is conjectured that the values of the Riemann zeta function at all positive integers are irrational.
In 1882, Lindemann proved that the number $\pi$ is transcendental (see, for example, [6, Theorem 1.2]). As a consequence, the even values $\zeta(2n)$ are irrational. In 1979, Apéry [1] famously demonstrated that the number $\zeta(3)$ is irrational. Subsequently, Beukers [3] provided an elementary proof of $\zeta(3)$ based on integral representations. In 2001, Keith Ball and Tanguy Rivoal [2, Théorème 1] applied Nesterenko’s criterion to prove that the $\mathbb{Q}$-vector space spanned by odd zeta values is infinite-dimensional:
$\operatorname{dim}_{\mathbb{Q}}\langle 1,\zeta(3),\zeta(5),\ldots,\zeta(2n+1)\rangle_{\mathbb{Q}}\geq\frac{1}{3}\log(2n+1).$
In the same year, Zudilin [7] demonstrated that at least one of the four numbers
$\zeta(5),\zeta(7),\zeta(9),\zeta(11)$
is irrational. Incidentally, the proof of the irrationality of $\zeta(3)$ can also be found in Zudilin’s recent book [8, Section 7].
To this day, it remains unknown whether $\zeta(5)\notin\mathbb{Q}$. It is also not known whether $\zeta(3)\notin\pi^{3}\mathbb{Q}$.
Recently, Shekhar Suman [5] claimed to have proved the irrationality of $\zeta(5)$. Unfortunately, the proposed proof is incorrect.
This note is organized as follows. In Section 2, we identify the logical flaws in the proof presented in [5]. In Section 3, we outline a standard approach to proving the irrationality of a number and highlight that the irrationality of $\zeta(2m+1)$ for general odd integers $12m+$ 1 remains an open and intriguing area of research.
2. Logical Fallacies in Suman’s Proof
In this section, we analyze the flaws in Shekhar Suman’s proof of [5, Theorem 1]. While the preliminary result, [5, Lemma 1], is correct and well-established, the proof of [5, Theorem 1] contains a critical logical error.
Suman claims that the linear Diophantine equation [5, Eq. (47)] has no integer solution, which he uses to derive the irrationality of $\zeta(5)$. However, we show that [5, Eq. (47)] indeed admits integer solutions, invalidating his conclusion.
Suman argues as follows:
Since $(d_{n},2d_{n})=d_{n}$, the above linear Diophantine equation has an integral solution if and only if $d_{n}\mid k_{i}b$. So we have
$d_{n}a-2d_{n}b=-k_{i}b\quad\text{where }d_{n}\mid k_{i}b,\quad 1\leq k_{i}\leq d_{n}-1,\quad n\geq b.$
He then claims that the following cases are impossible:
$d_{n}a-2d_{n}b=-k_{i}b\quad\text{where }d_{n}\mid k_{i}b,\quad 0\leq k_{i}\leq d_{n},\quad n\geq 1.$
He employs induction on $n$ to establish this claim. However, a closer inspection reveals a critical flaw in his argument at the base case $n=1$.
For $n=1$, the equation becomes:
$a-2b=-k_{i}b,\quad\text{where }0\leq k_{i}\leq 1\text{ and }1\mid k_{i}b.$
Solving this, Suman correctly concludes that the possible solutions are:
$a-2b=0\quad\text{or}\quad a-2b=-b.$
This implies that $a=2b$ or $a=b$, both of which are valid integer solutions. However, instead of acknowledging the existence of these integer solutions as a refutation of his claim, Suman proceeds to argue that these solutions lead to:
$\zeta(5)=\frac{a}{b}=2\quad\text{or}\quad\zeta(5)=1,$
which he deems "absurd" because it is well-known that $\zeta(5)$ is not an integer.
Here lies the critical error: the claim that equation [5, Eq. (48)] has no integer solutions is logically independent of whether $\zeta(5)$ is an integer or not. The well-established fact that $\zeta(5)$ is not an integer cannot be used to prove that [5, Eq. (48)] has no integer solutions. The existence of integer solutions for [5, Eq. (48)] at $n=1$ directly invalidates the induction base case and, consequently, the entire proof.
To summarize, the error in [5, Theorem 1] lies in conflating two unrelated facts: the solvability of the Diophantine equation [5, Eq. (48)] and the irrationality of $\zeta(5)$. The former is a purely algebraic property of the equation, while the latter is a number-theoretic property of the zeta function. By assuming $\zeta(5)$ is not an integer to argue against the solvability of [5, Eq. (48)], Suman undermines the logical foundation of his proof. As a result, Theorem 1 in [5] is incorrect.
In addition to the errors in the proof of [5, Theorem 1], similar logical fallacies can be identified in the proof of [5, Theorem 2], which aims to establish the irrationality of $\zeta(2m+1)$ for integers $m\geq 2$. While the specific details of the argument in [5, Theorem 2] differ from [5, Theorem 1], the underlying issue remains the same: the reliance on assumptions or conclusions that are independent of the solvability of the corresponding Diophantine equations.
3. A Standard Approach to Proving the Irrationality of a Number
In this section, we review a standard method for proving the irrationality of a number $\alpha$ (the criterion for irrationality), and show that Suman’s method does not meet the criterion.
Proposition 3.1.
Suppose that we can construct sequences of pairs of rational numbers $a_{n},b_{n}$ with the following properties:
  • (1) There is a small number $10<\varepsilon<$ 1 such that
    $0<\left|a_{n}\alpha-b_{n}\right|<\varepsilon^{n}$
    for all sufficiently large $n$.
  • (2) Let $d_{n}\in\mathbb{N}$ be the common denominator of $a_{n},b_{n}$:
    $d_{n}a_{n}\in\mathbb{Z},\quad d_{n}b_{n}\in\mathbb{Z}.$
    Assume that $d_{n}<D^{n}$ for some $D\in\mathbb{R}$.
  • (3) $D$ is not too big:
    $D\varepsilon<1.$
Then $\alpha$ is irrational.
Proof.
(by contradiction) Suppose that $\alpha$ is rational, $\alpha=\frac{p}{q}$ where $p,q\in\mathbb{Z},q>0$. Assumption (1) then becomes
$0<\left|a_{n}\frac{p}{q}-b_{n}\right|<\varepsilon^{n}\quad\text{for large }n.$
By multiplying through by $q$ and $d_{n}$, we obtain
$0<\left|d_{n}a_{n}p-d_{n}b_{n}q\right|<qd_{n}\varepsilon^{n}<qD^{n}\varepsilon^{n}.$
Since by assumption (3), $D\varepsilon<1$, the right-hand side tends to zero. Thus, we can find a large $n$ such that
$10<\left|\underbrace{(d_{n}a_{n})}_{\in\mathbb{Z}}p-\underbrace{(d_{n}b_{n})}_{\in\mathbb{Z}}q\right|<$ 1 .
But by (2), this is an integer between 0 and 1, which is a contradiction. ∎
Remark.
Besides Beukers [3], other interesting applications of Proposition 3.1 can be found in, for example, Huylebrouck [4].
Numerical verification shows that the $I_{n}$ in [5, Lemma 1] does not satisfy the condition (3) of Proposition 3.1. Therefore, $I_{n}$ cannot be used to prove the irrationality of $\zeta(5)$ via Proposition 3.1. In fact, mathematicians have been searching for a suitable integral representation of the form $a_{n}+b_{n}\zeta(5)$ that converges to 0, but such a representation has not yet been found. Therefore, the irrationality of $\zeta(2m+1)$ for $m\geq 2$ still remains an open problem worth exploring. Finally, finding a good integral representation for $\zeta(5)$ or even for $\zeta(2m+1)$ that can be used to approximate them continues to be a worthwhile pursuit.

### v3

@@
 where $B_{m}$ is the $m$-th Bernoulli number. It is conjectured that the values of the Riemann zeta function at all positive integers are irrational.
-In 1882, Lindemann proved that the number $\pi$ is transcendental (see, for example, [6, Theorem 1.2]).
+In 1882, Lindemann proved that the number $\pi$ is transcendental (see, for example, [7, Theorem 1.2]).
 As a consequence, the even values $\zeta(2n)$ are irrational.
@@
 $\zeta(5),\zeta(7),\zeta(9),\zeta(11)$
-is irrational. Incidentally, the proof of the irrationality of $\zeta(3)$ can also be found in Zudilin’s recent book [8, Section 7].
+is irrational. Incidentally, the proof of the irrationality of $\zeta(3)$ can also be found in Zudilin’s recent book [9, Section 7].
 To this day, it remains unknown whether $\zeta(5)\notin\mathbb{Q}$. It is also not known whether $\zeta(3)\notin\pi^{3}\mathbb{Q}$.
@@
 In Section 3, we outline a standard approach to proving the irrationality of a number and highlight that the irrationality of $\zeta(2m+1)$ for general odd integers $12m+$ 1 remains an open and intriguing area of research.
+In Section 4, we discuss recent advancements, including the groundbreaking work of Calegari, Dimitrov, and Tang [4], which introduces a powerful new method for proving irrationality and offers promising directions for future research.
 2. Logical Fallacies in Suman’s Proof
-In this section, we analyze the flaws in Shekhar Suman’s proof of [5, Theorem 1].
-While the preliminary result, [5, Lemma 1], is correct and well-established, the proof of [5, Theorem 1] contains a critical logical error.
-Suman claims that the linear Diophantine equation [5, Eq.
+In this section, we analyze the flaws in Shekhar Suman’s proof of [6, Theorem 1].
+While the preliminary result, [6, Lemma 1], is correct and well-established, the proof of [6, Theorem 1] contains a critical logical error.
+Suman claims that the linear Diophantine equation [6, Eq.
 (47)] has no integer solution, which he uses to derive the irrationality of $\zeta(5)$.
-However, we show that [5, Eq.
+However, we show that [6, Eq.
 (47)] indeed admits integer solutions, invalidating his conclusion.
@@
 which he deems "absurd" because it is well-known that $\zeta(5)$ is not an integer.
-Here lies the critical error: the claim that equation [5, Eq.
+Here lies the critical error: the claim that equation [6, Eq.
 (48)] has no integer solutions is logically independent of whether $\zeta(5)$ is an integer or not.
-The well-established fact that $\zeta(5)$ is not an integer cannot be used to prove that [5, Eq.
+The well-established fact that $\zeta(5)$ is not an integer cannot be used to prove that [6, Eq.
 (48)] has no integer solutions.
-The existence of integer solutions for [5, Eq.
+The existence of integer solutions for [6, Eq.
 (48)] at $n=1$ directly invalidates the induction base case and, consequently, the entire proof.
-To summarize, the error in [5, Theorem 1] lies in conflating two unrelated facts: the solvability of the Diophantine equation [5, Eq.
+To summarize, the error in [6, Theorem 1] lies in conflating two unrelated facts: the solvability of the Diophantine equation [6, Eq.
 (48)] and the irrationality of $\zeta(5)$.
 The former is a purely algebraic property of the equation, while the latter is a number-theoretic property of the zeta function.
-By assuming $\zeta(5)$ is not an integer to argue against the solvability of [5, Eq.
+By assuming $\zeta(5)$ is not an integer to argue against the solvability of [6, Eq.
 (48)], Suman undermines the logical foundation of his proof.
 As a result, Theorem 1 in [6] is incorrect.
-In addition to the errors in the proof of [5, Theorem 1], similar logical fallacies can be identified in the proof of [5, Theorem 2], which aims to establish the irrationality of $\zeta(2m+1)$ for integers $m\geq 2$.
-While the specific details of the argument in [5, Theorem 2] differ from [5, Theorem 1], the underlying issue remains the same: the reliance on assumptions or conclusions that are independent of the solvability of the corresponding Diophantine equations.
+In addition to the errors in the proof of [6, Theorem 1], similar logical fallacies can be identified in the proof of [6, Theorem 2], which aims to establish the irrationality of $\zeta(2m+1)$ for integers $m\geq 2$.
+While the specific details of the argument in [6, Theorem 2] differ from [6, Theorem 1], the underlying issue remains the same: the reliance on assumptions or conclusions that are independent of the solvability of the corresponding Diophantine equations.
 3. A Standard Approach to Proving the Irrationality of a Number
@@
 Besides Beukers [3], other interesting applications of Proposition 3.1 can be found in, for example, Huylebrouck [5].
-Numerical verification shows that the $I_{n}$ in [5, Lemma 1] does not satisfy the condition (3) of Proposition 3.1.
+Numerical verification shows that the $I_{n}$ in [6, Lemma 1] does not satisfy the condition (3) of Proposition 3.1.
 Therefore, $I_{n}$ cannot be used to prove the irrationality of $\zeta(5)$ via Proposition 3.1.
@@
 Finally, finding a good integral representation for $\zeta(5)$ or even for $\zeta(2m+1)$ that can be used to approximate them continues to be a worthwhile pursuit.
+4. Future Perspectives
+Recently, Calegari, Dimitrov, and Tang [4] made significant progress in the study of irrationality and $\mathbb{Q}$-linear independence of special values in number theory.
+Their work established several groundbreaking results, including the irrationality of the classical Dirichlet $L$-value
+$L(2,\chi_{-3})=\sum_{n=0}^{\infty}\left(\frac{1}{(3n+1)^{2}}-\frac{1}{(3n+2)^{2}}\right),$
+and demonstrated its linear independence with $1$ and $\pi^{2}$ over $\mathbb{Q}$.
+This result, formalized as [4, Theorem A] in their paper, represents a significant advancement in our understanding of special values of $L$-functions.
+Furthermore, their techniques generalized Zagier’s constructions and provided a new arithmetic holonomy bound framework.
+Their work also yielded [4, Corollary B], which proves the irrationality of numerous related series.
+For instance, they showed that certain linear combinations of $L(2,\chi_{-3})$, $\pi^{2}$ are irrational.
+These results expand the set of known irrational numbers derived from $L$-functions and highlight the potential of their methods for exploring similar problems.
+Moreover, [4, Theorem C] extends their analysis to logarithmic values, establishing the irrationality of $\log\left(1+\frac{1}{m}\right)\log\left(1+\frac{1}{n}\right)$ under the condition $|\frac{m}{n}-1|<\frac{1}{10^{6}}$ and proving the $\mathbb{Q}$-linear independence of the following four expressions:
+$1,\quad\log\left(1+\frac{1}{m}\right),\quad\log\left(1+\frac{1}{n}\right),\quad\log\left(1+\frac{1}{m}\right)\log\left(1+\frac{1}{n}\right).$
+These findings represent a major breakthrough in the study of special values of zeta and $L$-functions, with implications for broader problems in number theory, transcendence theory, and the study of modular forms and related automorphic objects.
+The techniques introduced by Calegari, Dimitrov, and Tang provide a robust foundation for future exploration, including the study of other special values and related constants in mathematics.